A lawn bench price at $523.89 was just discounted 10%. Then an additional 4% discount is added for paying cash.If you pay with cash, what will the bench cost you after the discounts are applied? B. If there was a 20% mark-up from the wholesale cost to get the price of $523.89, what was the wholesale cost or the cost to the dealer?
Answer by rfer(16322)

A)
523.89*0.90=471.50
471.50*0.96=$452.64
B)
523.89/1.20=$436.58